OpenStudy (anonymous):

Evaluate y = 3|x| – 3 for x = –1. (1 point) ?? Can u please help me? :'(

OpenStudy (aripotta):

just substitute -1 for x in the equation

OpenStudy (aripotta):

y = 3|-1| – 3

OpenStudy (aripotta):

then solve it

OpenStudy (anonymous):

so -3?? or 3

OpenStudy (aripotta):

y = 3|-1| – 3 y = 3(1) - 3 since the absolute value of -1 is 1 y = 3 - 3 y = 0 does that make sense?

OpenStudy (anonymous):

Oh, okay yeah it does now! Thank ouy so much! (:

OpenStudy (aripotta):

no problemo :]

OpenStudy (anonymous):

Okay, here one more Q i'm having problems with What is the value of the function when x = –2? y = x2 – 3x + 1 (1 point)

OpenStudy (anonymous):

and this is the last one

OpenStudy (aripotta):

is that x^2?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

–11 –10 10 11

OpenStudy (aripotta):

substitute -2 for x y = -2^2 – 3(-2) + 1

OpenStudy (anonymous):

-11??

OpenStudy (aripotta):

why do you think it's -11

OpenStudy (anonymous):

I don't knwo :(

OpenStudy (aripotta):

ok, do it in pieces what's -2^2?

OpenStudy (anonymous):

-8?

OpenStudy (aripotta):

no -2^2 = -2 times -2 = 4

OpenStudy (anonymous):

Ok

OpenStudy (aripotta):

and what's -3(-2)?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

+6 (positive 6)

OpenStudy (aripotta):

so y = -2^2 – 3(-2) + 1 y = 4 + 6 + 1 y = ?

OpenStudy (anonymous):

+11?

OpenStudy (anonymous):

11

OpenStudy (aripotta):

yes :]

OpenStudy (anonymous):

YAY!! :) Thanks so much

OpenStudy (aripotta):

mhmm no problem

OpenStudy (anonymous):

Thanks for not just giving me the answer i like to be helped not given the answer(:

OpenStudy (aripotta):

well i hope it all made sense to you

OpenStudy (anonymous):

Ya know? So i can jearn how to do it and you made that possible for me tyhank you sooo much again

OpenStudy (anonymous):

It did a little bit mixed up but thats just my dyslexic acting up

OpenStudy (anonymous):

IDyslexic is hard but not impossible to work with

OpenStudy (anonymous):

I'm homeschooled because every teacher i had told je i was dumb, or got frusterated and didn't help me

OpenStudy (aripotta):

i couldn't really tell you were dyslexic, i thought it was just your name lol i don't think you're dumb. i think you can really do this

OpenStudy (anonymous):

Thanks(: maybe i should change my name lo.l

OpenStudy (aripotta):

you can't change your name, you're stuck with it forever mwuhaha

OpenStudy (anonymous):

lol